HUBBLE SPACE TELESCOPE DAILY REPORT #3092 PERIOD COVERED: 0000Z (UTC) 04/10/02 - 0000Z (UTC) 04/11/02 Daily Status Report as of 101/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.2 Completed ACS/HRC/WFC 9292 (The Nature of Galaxies at z > 4) The Advanced Camera for Surveys (HRC and WFC) was used to further look into recent discoveries of a number of galaxies and quasars at redshifts greater than 5 that has identified the z>5-6 epoch as key to understanding the earliest formation phases for galaxies. However, establishing the characteristics and properties of these earliest galaxies is proving to be a particularly difficult. They are faint, with I{AB} magnitudes around 26-27. Thus, substantial investments of time are needed to obtain high S/N images, while ground-based spectroscopy, even with 8-10 m class telescopes, has provided little more than redshifts. Establishing the physical properties of these galaxies will be a challenge for the foreseeable future. However, there is a subset of this high redshift population that is amenable to more detailed study. These are sources that have been strongly lensed by low redshift clusters. No problems were reported. 1.3 Completed ACS/HRC 9032 (CCD Flash Calibration) The Advanced Camera for Surveys (HRC) was used to provide a preliminary baseline set of CCD FLASH exposure reference images for each current level/shutter-side/detector combination, for the pair of FLASH LEDs on the instrument side currently in use. It also tested the short-term repeatability at the shortest FLASH exposure times that are expected to be used {1.0 sec}. The status buffer message, mentioned in HSTAR 8601 and yesterday's reported, occurred during this proposal. Per subsequent information from SE and the STScI, this error did not affect science or image quality, and may occur twice more in the current SMS.. There were no other reported problems. 1.4 Completed Two Sets of STIS/CCD 9070 (A Census of Nuclear Star Clusters in Late-Type Spiral Galaxies: II. Spectroscopy and Stellar Populations) The Space Telescope Imaging Spectrograph (CCD) was used to investigate spiral galaxies that have a prominent star cluster in their dynamical center. Statistics for cluster frequency, size, and luminosity remain incomplete. The proposal completed as planned. 1.5 Completed Twenty-one Sets of ACS/WFC 8948 (CTE Test) The Advanced Camera for Surveys (WFC) was used to measure the baseline charge transfer efficiency {CTE} performance of the WFC detectors, emulating the CTE ground calibration testing. Also, a similar program will perform periodic measurements in order to track the change in the CTE as damage due to radiation exposure accumulates. All observations were successful. 1.6 Completed STIS/CCD 9126 (Galaxy Mass and the Fate of Luminous, Blue Compact Galaxies at z~0.6) The Space Telescope Imaging Spectrograph (CCD) was used to obtain long-slit spectra for a sample of 6 luminous, blue, compact galaxies {LBCGs} at z ~ 0.6. Despite being very luminous, LBCGs have velocity widths Sigma ~ 60 kms and half-light radii r_e ~ 0.5'' {or R_e ~ 3Kpc}. Small sizes and velocity widths suggest LBCGs are low- mass stellar systems , while their blue colors, strong emission lines and low M/L-ratios indicate they are undergoing a major starburst. I f the star-forming process halts after the current burst, models predict that LBCGs will fade by ~2-4 magnitudes after a few Gyrs to reach the low luminosities and surface brightnesses characteristic of spheroidal galaxies. The proposal completed with no reported problems. 1.7 Completed Two Sets of WF/PC-2 9145 (A Snapshot Survey of the Optically Selected Type-2 Quasars) The WF/PC-2 was used to observe an identified population of emission-line objects in DPOSS, which can be plausibly interpreted as the long-sought type-2 quasars. They have high-ionization Seyfert-2 like spectra, but with narrow-line luminosities comparable to those of the luminous type-1 quasars in the same redshift range. This population may be a major contributor to the cosmic hard x-ray background. It is proposed to obtain multi-color images of a representative sample of these objects, in order to examine their morphology. We may be able to detect point-like nuclei which are not detectable in ground-based images, the dust lanes hiding them from our view, possible evidence for tidal interactions and the overall morphology of their hosts, etc. The proposal completed with no reported problems. 1.8 Completed STIS/CCD 9131 (Imaging the Host Galaxies of High Redshift Type Ia Supernovae) The Space Telescope Imaging Spectrograph (CCD) was used to complete the snapshot survey of distant galaxies of known redshift which hosted supernovae {SNe} of Type Ia found via the Supernova Cosmology Project {SCP}. No problems were seen. 1.9 Completed STIS/CCD 9139 (Variability in the UV Spectrum of 3C 279: Testing Models for the Gamma-Ray Emission in Blazars) The Space Telescope Imaging Spectrograph (CCD) was used to determine whether gamma- rays are produced on the scale of the broad line region, using multi-epoch UV snapshots of 3C 279 to find and characterize its LyAlpha variability. This will validate the mirror model for generating the intense gamma-ray emission from blazars. 3C 279 is the best-studied blazar from radio to gamma-ray wavelengths, showing frequent large flares over days to months. Variability in the UV continuum and LyAlpha will directly reveal any coupling between the jet ionizing flux and the broad-line region emission, providing clues to the physics and energetics of all radio-loud AGN. All observations completed without reported incident. 1.10 Completed Two Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 9 Successful: 9 Scheduled Re-acquisitions: 8 Successful: 8 2.2 FHST Updates: Scheduled: 19 Successful: 18 The update at 100/204658Z failed 2.3 Operations Notes: The NICMOS Cryo Cooler cool down continues nominally. As of yesterday morning, the NICMOS Outlet Temperature (MNPNCOLT) was 80.92K (expected end point 76K), the NCC TA Inlet Temperature (MNPTAILT) was 68.69K (expected end point 68K), the CLI Temperature (MNPLINTT) was 71.53K (expected end point 71K), and the NICMOS Inlet Temperature (MNRNCILT) was 64.07K (expected endpoint 63K). During the day, the following operations notes were utilized: Adjust ACS error count limit (twice) ACS STS 935 Change limits for MCE-2 reset detection Using ROP DF-18A, the engineering status buffer limit for SESBSLD was updated twice. The STIS MCE-2 reset at 100/180330Z while the low voltage was on and while outside any SAA interval. MAMA-2 will be recovered at the next high voltage event by normal SMS commanding. Per an operations request, the "wait for reservoir" feature of the CPL control law was disabled at 100/1806Z. Per HSTAR 8603, a 486 engineering status buffer message at 101/004415Z indicated that the FGS sequential attitude update failed because the error was too large to correct. Further analysis awaits the receipt of SSR engineering data. JCLM5V (ACS Cal Lamp -5V Power) flagged out-of-limits intermittently beginning a 101/075114Z and ending at 101/082814Z. HSTAR 8602 was written. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW